OpenStudy (anonymous):

differentiate wrt x y=cos^-1((a-bsinx)/(a+bsinx))

myininaya (myininaya):

implicitly find y' from the equation: \[\cos(y)=\frac{a-bsin(x)}{a+bsin(x)}\]

OpenStudy (anonymous):

can u tell me the first step..

OpenStudy (anonymous):

Take the derivative of both sides: \[\frac{d}{dx}\cos(y)=\frac{d}{dx}\left[\frac{a-b\sin x}{a+b\sin x}\right]\] Chain rule for the left, quotient for the right: \[-\sin y\frac{dy}{dx}=\frac{(a+b\sin x)\dfrac{d}{dx}[a-b\sin x]-(a-b\sin x)\dfrac{d}{dx}[a+b\sin x]}{(a+b\sin x)^2}\] Some derivatives: \[-\sin y\frac{dy}{dx}=\frac{(a+b\sin x)(-b\cos x)-(a-b\sin x)(b\cos x)}{(a+b\sin x)^2}\] Take out a common factor and simplify what you can \[\begin{align*}-\sin y\frac{dy}{dx}&=\frac{-b\cos x\bigg(a+b\sin x+a-b\sin x\bigg)}{(a+b\sin x)^2}\\ \sin y\frac{dy}{dx}&=\frac{2ab\cos x}{(a+b\sin x)^2} \end{align*}\] and that's most of the work. Solve for \(\dfrac{dy}{dx}\).

OpenStudy (anonymous):

\[y=\cos^{-1} \frac{ a-b \sin x }{ a+b \sin x }\] \[\frac{ dy }{ dx }=-\frac{ 1 }{ \sqrt{1-\left( \frac{ a-b \sin x }{ a+b \sin x } \right)^2} }\frac{ d }{ dx }\left( \frac{ a-b \sin x }{ a+b \sin x } \right)\] \[=-\frac{ 1 }{ \frac{\sqrt{a^2+b^2sin^2x+2 ab \sin x-(a^2+b^2sin^2x-2 ab \sin x)} }{ a+b \sin x } }*\frac{ \left( a+b \sin x \right)\left(- b \cos x \right)-\left( a-b \sin x \right)\left( b \cos x \right) }{ \left( a+b \sin x \right)^2 }\] \[=\frac{ -a b \cos x-b^2\sin x \cos x-a b \cos x+b^2\sin x \cos x }{ \left( a+b \sin x \right)\sqrt{4 ab \sin x} }\] \[=-\frac{ -2 ab \cos x }{ 2\sqrt{a b \sin x} \left( a+b \sin x \right)}=?\]